OpenStudy (anonymous):

1. use the distributive property to simplify the expression . ( -1 ) (4-c ) ( 1point ) A . 4-c b. -4 + c c. 4 + c d . -4 -c

OpenStudy (sleepyhead314):

distributive property is a(b-c) = a*b - a*c so (-1)(4 - c) becomes (-1)*(4) - (-1)(c) = ?

OpenStudy (triciaal):

to use the distributive property multiply each term inside the parenthesis by what is outside ( -1 ) (4-c ) =(-1)*(4) +(-1)*(-c)

OpenStudy (anonymous):

so which one is it ? I'm a lil confused

OpenStudy (sleepyhead314):

it will get you the same answer either way just remember that negative times negative = positive